# reference.bib

### v2

@@ Abstract
3D lane detection is essential in autonomous driving as it extracts structural and traffic information from the road in three-dimensional space, aiding self-driving cars in logical, safe, and comfortable path planning and motion control. Given the cost of sensors and the advantages of visual data in color information, 3D lane detection based on monocular vision is an important research direction in the realm of autonomous driving, increasingly gaining attention in both industry and academia. Regrettably, recent advancements in visual perception seem inadequate for the development of fully reliable 3D lane detection algorithms, which also hampers the progress of vision-based fully autonomous vehicles. We believe that there is still considerable room for improvement in 3D lane detection algorithms for autonomous vehicles using visual sensors, and significant enhancements are needed. This review looks back and analyzes the current state of achievements in the field of 3D lane detection research. It covers all current monocular-based 3D lane detection processes, discusses the performance of these cutting-edge algorithms, analyzes the time complexity of various algorithms, and highlights the main achievements and limitations of ongoing research efforts. The survey also includes a comprehensive discussion of available 3D lane detection datasets and the challenges that researchers face but have not yet resolved. Finally, our work outlines future research directions and invites researchers and practitioners to join this exciting field.
Figure 1: Chronological overview of the most prestigious monocular image-based 3D lane detection. For methods without abbreviations, we use circled numbers as substitutes, and the corresponding paper names are listed above the timeline chart.
@@ I INTRODUCTION
In recent advancements fueled by artificial intelligence, autonomous driving technology has seen rapid evolution, progressively reshaping the paradigms of human transportation. Equipped with an array of sensors, autonomous vehicles emulate human sensory capabilities such as vision and hearing to perceive their surroundings and interpret traffic scenarios for safe navigation[wei2021multi]. Key among these sensors are LiDAR, high-resolution cameras, millimeter-wave radars, and ultrasonic radars[zou2021comparative, ma2022automatic], which facilitate feature extraction and object classification combined with high-precision mapping to discern obstacles and the vehicular traffic landscape[zhao2024curbnet]. Visual sensors, the most extensively utilized in autonomous vehicles, serve as the primary means for environmental perception, encompassing lane detection, traffic light analysis, road sign detection and recognition, vehicle tracking, pedestrian detection, and short-term traffic prediction[chen2021data]. The processing and understanding of visual scenes in autonomous driving, which includes analysis of traffic lights, identification of traffic signs, lane detection, and the detection of nearby pedestrians and vehicles, underpin more robust and safer commands for maneuvers such as steering, overtaking, lane changes, or braking[usman2020speed, ma2024dataset]. This integration of sensory data and environmental understanding seamlessly transitions into the domain of scene understanding in autonomous driving, which is pivotal for advancing vehicular autonomy and ensuring road safety.
Scene understanding represents one of the most challenging aspects within the domain of autonomous driving[ma2023self]. Without the capability for comprehensive scene understanding, navigating an autonomous vehicle safely through traffic lanes can be as daunting as navigating the world blindfolded for humans. Lane detection, in particular, stands out as a pivotal and challenging task within the realm of scene comprehension[xu2023centerlinedet]. Lanes serve as the most common traffic elements on roads, acting as crucial markers that segment the roadways to ensure the safe and efficient passage of vehicles[tang2021review]. Lane detection technology, which automatically identifies road markings, is indispensable; autonomous vehicles lacking this capability could lead to traffic congestion and even severe collisions, thereby compromising passenger safety. Thus, lane detection plays a critical role in the ecosystem of autonomous driving. Unlike typical objects, lane markings occupy only a slender portion of the road scene and are widely distributed, making them uniquely challenging to detect[sun2019accurate]. The task is further complicated by a variety of lane markings, insufficient lighting, obstructions, and interference from similar textures, which are common in many driving scenarios, thus exacerbating the intrinsic challenges associated with lane detection.
@@ I-A Background and Related Works
Due to advancements in deep learning technologies, researchers have developed numerous strategies that have significantly simplified, expedited, and enhanced the task of lane detection[tabelini2021keep]. Concurrently, as deep learning becomes increasingly widespread and as new concepts continuously emerge[ma20233d], the domain of lane detection has seen a further specialization and refinement in its approaches. Reflecting on the mainstream research trajectories within this domain, camera-based lane detection methods can be principally segmented into two-dimensional (2D) and three-dimensional (3D) lane detection paradigms[chen2022persformer].
@@
2D lane detection. [jin2022eigenlanes, liu2021end] methods aim to accurately delineate the shape and position of lanes within images, primarily employing four distinct approaches: segmentation-based, anchor-based, keypoint-based, and curve-based strategies. (1) Segmentation-based approaches frame 2D lane detection as a pixel-level classification challenge, generating lane masks[zheng2021resa, neven2018towards, pan2018spatial, zheng2021resa]. These methods cultivate lane groupings through the exploration of effective semantic features and subsequent post-processing, although they are marked by their high computational costs. (2) Anchor-based approaches are lauded in 2D lane detection for their simplicity and efficiency, typically utilizing linear anchor points to regress positional offsets relative to targets [li2019line, tabelini2021keep, zheng2022clrnet]. To circumvent the constraints of linear anchors, [jin2022eigenlanes]employs an intrinsic lane space generating various lane candidates. Heuristically designed row anchors classify row pixels as lanes [yoo2020end, liu2021condlanenet], further evolved into mixed (row and column) anchors in [qin2022ultra] to alleviate positioning errors for side lanes. Moreover, this method significantly enhances inference speed. (3) Keypoint-based approaches offer a more flexible and sparse modeling of lane positions[ko2021key, qu2021focus, wang2022keypoint], initiating with the estimation of point locations, followed by associating keypoints belonging to the same lane using varying schemes. The primary strategy in [su2021structure, tabelini2021keep] is to predict 2D lanes by representing lane structures with predefined keypoints and regressing offsets between sampled and predefined points. Despite promising results, these methods lack the flexibility to adapt to complex lane configurations due to their fixed point design. (4) Curve-based approaches [tabelini2004lane, van2019end, liu2021end]focus on fitting lane lines through various curve equations and specific parameters, transforming 2D lane detection into a curve parameter regression challenge by detecting start and end points alongside curve parameters. Although there have been promising advancements in 2D lane detection, a significant gap remains between the 2D outcomes and the real-world application requirements, notably the precise 3D positioning.
@@
3D lane detection. Due to the inherent lack of depth information in 2D lane detection, projecting these detections into 3D space can result in inaccuracies and reduced robustness[bai2024curveformer++]. Consequently, numerous researchers have shifted their focus towards lane detection within the 3D domain. Deep learning-based 3D lane detection methodologies primarily bifurcate into CNN-based and Transformer-based approaches, initially constructing dense Bird’s Eye View (BEV) feature maps, followed by the extraction of 3D lane information from these intermediate representations[huang2023anchor3dlane].
@@
CNN-based methods predominantly include D-LaneNet [garnett20193d], which proposes a dual-path architecture utilizing Inverse Perspective Mapping (IPM) to transpose features and detect lanes through vertical anchor regression. 3D-LaneNet+ [efrat20203d] segments BEV features into non-overlapping units, addressing the limitations of anchor orientation through lateral offsets, angles, and elevation shifts relative to the unit’s center. GenLaneNet [guo2020gen] pioneers the use of a fictitious top-view coordinate system for better feature alignment and introduces a two-stage framework to decouple lane segmentation and geometric encoding. BEVLaneDet [wang2023bev] employs a virtual camera to ensure spatial consistency and adapts to more complex scenarios through a keypoint-based representation of 3D lanes. GroupLane [li2023grouplane] innovates within the BEV by introducing a row-based classification method, which accommodates lanes in any direction and interacts with feature information within instance groups.
@@
Transformer-based methods include CLGo [liu2022learning], presenting a two-stage framework capable of estimating camera posture from the image and lane decoders based on BEV features. PersFormer [chen2022persformer] constructs dense BEV queries using offline camera poses, unifying 2D and 3D lane detection under a Transformer-based framework. STLanes3D [wang2022spatio] predicts 3D lanes using fused BEV features and introduces the 3DLane-IOU loss to couple lateral and height errors. Anchor3DLane [huang2023anchor3dlane], a CNN-based method, regresses 3D lanes directly from image features based on 3D anchors, significantly reducing computational overhead. CurveFormer [bai2022curveformer] leverages sparse query representations and cross-attention mechanisms within Transformers to effectively regress polynomial coefficients of 3D lanes. LATR [luo2023latr], building on CurveFormer’s query anchor modeling, constructs lane-aware query generators and dynamic 3D ground position embeddings. CurveFormer++[bai2024curveformer++] proposes a single-stage Transformer detection method that does not require image feature view transformation and directly infers 3D lane detection results from perspective image features.
@@ I-B Challenges and Motivation
Accurately estimating the 3D position of lane markings based on monocular vision presents numerous challenges. Firstly, real-world data for 3D lane detection is highly variable due to the diversity of weather [kuang2020real], road types, markings, and environmental conditions, which complicates the training of models that can generalize well across different scenarios [campbell2010autonomous, zhang2023perception]. Furthermore, processing 3D data for lane detection requires substantial computational resources, which is particularly critical in autonomous driving applications where low latency is imperative. Reducing the processing time of algorithms without compromising their performance is also a significant challenge in such real-time applications. Recently, the community lacks a consolidated, unified, single point of reference for ascertaining the current level of maturity of camera-based 3D lane detection techniques for autonomous driving.
Considering the aforementioned challenges and the importance of vision sensor-based semantic segmentation in accurate scene understanding and parsing, we have accumulated existing research contributions and outcomes in this survey. The main research questions highlighted in this survey are as follows: 1) Do existing datasets possess the potential for 3D lane detection in complex visual scenes? 2) What are the model sizes and inference speeds of current methods, and can these methods meet the real-time requirements of autonomous vehicles? 3) Can current methods effectively perform 3D lane detection in complex visual scenes containing uncertainties such as fog and rain?
@@ I-C Contributions
    A comprehensive introduction to 3D lane detection techniques, defining the generic pipeline and explaining each step individually. This helps newcomers in the field to quickly grasp prior knowledge and contributions from previous research, especially in the context of autonomous driving. To the best of our knowledge, this is the first survey on camera-based 3D lane detection.
@@ I-D Review Methodology
The research works discussed in this survey were retrieved using different keywords such as 3D lane detection in autonomous vehicles, vision-based 3D lane detection, and learning-based 3D lane detection. Most of the retrieved papers are directly relevant to the research topic, but there are also some exceptions, such as multi-modal methods [luo2022m, bai2018deep], and point cloud-based approaches [jung2018real], which are less related to the theme of this survey. Additionally, the aforementioned keywords were searched in multiple repositories, including Web of Science and Google Scholar, to ensure the retrieval of relevant content. Inclusion criteria ensured that a paper was recognized by experts in autonomous driving, based on factors such as citation count or the impact of prior work. It is worth mentioning that when consulting the literature, no work on monocular 3D lane detection based on traditional methods was found. This may be because, unlike single-camera 2D lane detection, which only requires identifying pixels belonging to the lane in a 2D image, monocular 3D lane detection requires determining the three-dimensional position information of the lane in the 3D space using the 2D image. Without the assistance of distance measuring sensors such as LiDAR [^1] https://en.wikipedia.org/wiki/Lidar or prediction through deep learning, it is difficult to achieve.
@@
The remainder of this paper is structured as follows: Section II provides an overview of related works for 3D lane detection in autonomous driving. Section III introduces performance evaluation of 3D lane detection. In section IV, commonly used datasets are introduced. And, Next, in section V, we introduce the current challenges and prospects for future 3D lane detection. Finally, Section VI concludes the paper.
@@
Figure 2: An overview of the 3D-LaneNet architecture. Information is processed in two parallel streams or pathways: the image-view pathway and the top-view pathway. This is called the dual-pathway backbone.
II MONOCULDAR 3D LANE DETECTION IN AD
With the rapid development of deep learning and autonomous driving, deep learning-based monocular lane detection has received increasing attention from both the industry and academia. In the field of monocular lane detection, early work mainly focused on 2D lane detection. As autonomous driving technology matures and the demand for cost reduction increases, higher requirements are placed on lane detection, aiming to predict the three-dimensional information of lane lines from a single image.Therefore, starting from 2018, there have been successive works on monocular 3D lane detection. As shown in Fig. 1, the figure provides a chronological overview of the monocular 3D lane detection algorithms. It can be seen that with the passage of time, more and more works have emerged, indicating that this field is becoming increasingly popular. In this figure, the green arrows represent CNN-based methods, and the orange arrows represent Transformer-based methods.
Among these methods, 3D-LaneNet [garnett20193d] is a pioneering work in the field of monocular 3D lane detection. 3D-LaneNet introduces a network that can directly predict three-dimensional lane information in road scenes from monocular images. This work is the first to solve the task of three-dimensional lane detection using on-board monocular vision sensor. 3D-LaneNet introduces two novel concepts: intra-network feature map inverse perspective mapping (IPM) and anchor-based lane representation. The intra-network IPM projection facilitates dual representation information flow in both front view and bird’s-eye view. The anchor-based lane output representation supports an end-to-end training approach, which is different from common heuristic methods that equate the problem of detecting three-dimensional lane lines to object detection problems. The overview of 3D-LaneNet is depicted in the Fig. 2. Inspired by works such as FCOS [tian2019fcos] and CenterNet [zhou2019objects], 3D LaneNet+ [efrat20203d] is an anchor-free 3D lane detection algorithm that can detect 3D lane lines of arbitrary topological structures. The authors of 3D LaneNet+ followed the dual-stream network of 3D LaneNet that separately processes the image view and bird’s-eye view and extended it to support the detection of 3D lane lines with more topological structures. Instead of predicting the entire lane as a whole, 3D LaneNet+ detects small lane segments that lie within the cell and their attributes (position, orientation, height). In addition, the method learns a global embedding for each cell to cluster small lane segments into complete 3D lane information. In [jiang2022robust], the Jiang et al. designed a two-stage 3D lane detection network where each stage is trained separately. The first subnetwork focuses on lane image segmentation, while the second subnetwork focuses on predicting the 3D lane structure based on the segmentation outputs from the first subnetwork. The Efficient Channel Attention (ECA) [wang2020eca] attention mechanism and the Convolutional Block Attention Module (CBAM) [woo2018cbam] attention mechanism were introduced in each stage, respectively, resulting in improved segmentation performance and accuracy in 3D lane detection. In [guo2020gen], Guo et al. proposed Gen-LaneNet, a general and scalable 3D lane detection method for detecting 3D lane lines from a single image, as shown in Fig. 3. The authors introduced a novel geometry-guided lane anchor representation and applied specific geometric transformations to directly compute the real 3D lane points from the network output. This anchor design extends intuitively from the anchor design in 3D-LaneNet. This method aligns the anchor coordinates with the underlying bird’s-eye-view features, enhancing its robustness in handling unfamiliar scenes. Additionally, this paper presents a scalable two-stage framework enabling independent learning of the image segmentation subnetwork and the geometry encoding subnetwork, thus significantly reducing the number of 3D labels required for training. Furthermore, this paper introduced a highly realistic synthetic image dataset with rich visual variations for developing and evaluating 3D lane detection methods.

### v3

@@ Abstract
3D lane detection is essential in AD as it extracts structural and traffic information from the road in 3D space, aiding autonomous vehicles in logical, safe, and comfortable path planning and motion control. Given the cost of sensors and the advantages of visual data in color information, 3D lane detection based on monocular vision is an important research direction in the realm of AD that increasingly gains attention in both industry and academia. Nevertheless, recent advancements in visual perception seem inadequate for the development of fully reliable 3D lane detection algorithms, which also hampers the progress of vision-based fully autonomous vehicles. We believe that it still leaves an open and interesting problem for improvement in 3D lane detection algorithms for autonomous vehicles using visual sensors, and significant enhancements are essentially required. This review summarizes and analyzes the current state of achievements in the field of 3D lane detection research. It covers all current monocular-based 3D lane detection processes, discusses the performance of these cutting-edge algorithms, analyzes the time complexity of various algorithms, and highlights the main achievements and limitations of ongoing research efforts. The survey also includes a comprehensive discussion of available 3D lane detection datasets and the challenges that researchers encounter but have not yet resolved. Finally, our work outlines future research directions and invites researchers and practitioners to join this exciting field.
Figure 1: Chronological overview of 3D lane detection based on monocular images. Green arrows represent methods based on CNNs, orange arrows represent methods based on Transformers, and blue arrows represent hybrid architecture methods.
@@ I INTRODUCTION
In recent advancements fueled by artificial intelligence, AD technology has seen rapid evolution, progressively reshaping the paradigms of human transportation. Equipped with an array of sensors, autonomous vehicles emulate human sensory capabilities such as vision and hearing to perceive their surroundings and interpret traffic scenarios for safe navigation [wei2021multi]. Key among these sensors are LiDAR, high-resolution cameras, millimeter-wave radars, and ultrasonic radars [zou2021comparative, ma2022automatic], which facilitate feature extraction and object classification combined with high-precision mapping to discern obstacles and the vehicular traffic landscape [zhao2024curbnet]. Visual sensors, which are most extensively utilized in autonomous vehicles, serve as the primary means for environmental perception, encompassing lane detection, traffic light analysis, road sign detection and recognition, vehicle tracking, pedestrian detection, and short-term traffic prediction [chen2021data]. The processing and understanding of visual scenes in AD, which includes analysis of traffic lights, identification of traffic signs, lane detection, and the detection of nearby pedestrians and vehicles, underpin more robust and safer commands for maneuvers such as steering, overtaking, lane changes, or braking [usman2020speed, ma2024dataset]. This integration of sensory data and environmental understanding seamlessly transitions into the domain of scene understanding in AD, which is pivotal for advancing vehicular autonomy and ensuring road safety.
Scene understanding represents one of the most challenging aspects within the domain of AD [ma2023self]. Without the capability for comprehensive scene understanding, navigating an autonomous vehicle safely through traffic lanes can be as daunting as navigating the world blindfolded for humans. Lane detection, in particular, stands out as a pivotal and challenging task within the realm of scene comprehension [xu2023centerlinedet]. Lanes serve as the most common traffic elements on roads, acting as crucial markers that segment the roadways to ensure the safe and efficient passage of vehicles [tang2021review]. Lane detection technology, which automatically identifies road markings, is indispensable; autonomous vehicles lacking this capability could lead to traffic congestion and even severe collisions, thereby compromising passenger safety. Thus, lane detection plays a critical role in the ecosystem of AD. Unlike typical objects, lane markings occupy only a slender portion of the road scene and are widely distributed, making them uniquely challenging to detect [sun2019accurate]. The task is further complicated by a variety of lane markings, insufficient lighting, obstructions, and interference from similar textures, which are common in many driving scenarios, thus exacerbating the intrinsic challenges associated with lane detection.
@@ I-A Background and Related Works
Due to advancements in deep learning technologies, researchers have developed numerous strategies that have significantly simplified, expedited, and enhanced the task of lane detection [tabelini2021keep]. Concurrently, as deep learning becomes increasingly widespread and as new concepts continuously emerge [ma20233d], the domain of lane detection has seen a further specialization and refinement in its approaches. Reflecting on the mainstream research trajectories within this domain, camera-based lane detection methods can be principally segmented into 2D and 3D lane detection paradigms [chen2022persformer].
@@
2D lane detection. In [jin2022eigenlanes, liu2021end], these works aim to accurately delineate the shape and position of lanes within images, primarily employing four distinct approaches: segmentation-based, anchor-based, keypoint-based, and curve-based strategies. (1) Segmentation-based approaches frame 2D lane detection as a pixel-level classification challenge, generating lane masks [zheng2021resa, neven2018towards, pan2018spatial, zheng2021resa]. These methods cultivate lane groupings through the exploration of effective semantic features and subsequent post-processing, although they are marked by their high computational costs. (2) Anchor-based approaches are lauded in 2D lane detection for their simplicity and efficiency, typically utilizing linear anchor points to regress positional offsets relative to targets [li2019line, tabelini2021keep, zheng2022clrnet]. To circumvent the constraints of linear anchors, various lane candidates are generated by an intrinsic lane space in [jin2022eigenlanes]. Heuristically designed row anchors classify row pixels as lanes [yoo2020end, liu2021condlanenet], further evolved into mixed (row and column) anchors in [qin2022ultra] to alleviate positioning errors for side lanes. Moreover, this method significantly enhances inference speed. (3) Keypoint-based approaches offer a more flexible and sparse modeling of lane positions [ko2021key, qu2021focus, wang2022keypoint], initiating with the estimation of point locations, followed by associating keypoints belonging to the same lane using varying schemes. The primary strategy in [su2021structure, tabelini2021keep] is to predict 2D lanes by representing lane structures with predefined keypoints and regressing offsets between sampled and predefined points. Despite promising results, these methods lack the flexibility to adapt to complex lane configurations due to their fixed point design. (4) Curve-based approaches [tabelini2021polylanenet, van2019end, liu2021end]focus on fitting lane lines through various curve equations and specific parameters, transforming 2D lane detection into a curve parameter regression challenge by detecting start and end points alongside curve parameters. Although there have been promising advancements in 2D lane detection, a significant gap remains between the 2D outcomes and the real-world application requirements, notably the precise 3D positioning.
@@
3D lane detection. Due to the inherent lack of depth information in 2D lane detection, projecting these detections into 3D space can result in inaccuracies and reduced robustness [bai2024curveformer++]. Consequently, numerous researchers have shifted their focus towards lane detection within the 3D domain. Deep learning-based 3D lane detection methodologies primarily bifurcate into CNN-based methods, Transformer-based methods, and hybrid methods.
@@
Representative CNN-based methods predominantly include 3D-LaneNet [garnett20193d], which proposes a dual-path architecture utilizing Inverse Perspective Mapping (IPM) to transpose features and detect lanes through vertical anchor regression. 3D-LaneNet+ [efrat20203d] segments bird’s eye view (BEV) features into non-overlapping units, addressing the limitations of anchor orientation through lateral offsets, angles, and elevation shifts relative to the unit’s center. GenLaneNet [guo2020gen] pioneers the use of a fictitious top-view coordinate system for better feature alignment and introduces a two-stage framework to decouple lane segmentation and geometric encoding. BEVLaneDet [wang2023bev] employs a virtual camera to ensure spatial consistency and adapts to more complex scenarios through a keypoint-based representation of 3D lanes. Anchor3DLane [huang2023anchor3dlane] regresses 3D lanes directly from image features based on 3D anchors, significantly reducing computational overhead. GroupLane [li2023grouplane] innovates within the BEV by introducing a row-based classification method, which accommodates lanes in any direction and interacts with feature information within instance groups. DecoupleLane [han2023decoupling] divides 3D lane detection into two parts: curve modeling and ground height regression. By representing lanes in the BEV space and modeling ground height separately, this method effectively addresses the fluctuations caused by uneven road surfaces.
@@
Transformer-based methods include PersFormer [chen2022persformer], which constructs dense BEV queries using offline camera poses, unifying 2D and 3D lane detection under a Transformer-based framework. CurveFormer [bai2022curveformer] leverages sparse query representations and cross-attention mechanisms within Transformers to effectively regress polynomial coefficients of 3D lanes. In [chen2023efficient], Chen et al. introduce an efficient transformer for 3D lane detection that jointly learns BEV and lane representations. Unlike methods using IPM for BEV transformation before lane detection, the model uses a decomposed cross-attention mechanism to learn both representations with supervision. LATR [luo2023latr], building on CurveFormer’s query anchor modeling, constructs lane-aware query generators and dynamic 3D ground position embeddings. CurveFormer++[bai2024curveformer++] proposes a single-stage Transformer detection method that does not require image feature view transformation and directly infers 3D lane detection results from perspective image features.
A series of methods also utilize hybrid network architectures, combining CNN and Transformer for 3D lane detection. CLGo [liu2022learning] proposes a two-stage framework. The first stage uses a CNN backbone for feature extraction and a transformer-encoder that aggregates nonlocal relations among spatial features. Additionally, it estimates the camera pose by introducing an auxiliary 3D lane task and geometric constraints. In the second stage, the estimated camera pose is used to generate a BEV image, which is then used to predict 3D lanes. Anchor3DLane [huang2023anchor3dlane] uses a CNN backbone and Transformer layer to extract features and directly predict 3D lanes from front-view images, bypassing the need for inverse perspective mapping. It defines 3D lane anchors as rays and projects them onto the feature map to capture necessary structural and contextual information for accurate 3D lane detection.
@@ I-B Challenges and Motivation
Accurately estimating the 3D position of lane markings based on monocular vision presents numerous challenges. Firstly, real-world data for 3D lane detection is highly variable due to the diversity of weather [kuang2020real], road types, markings, and environmental conditions, which complicates the training of models that can generalize well across different scenarios [campbell2010autonomous, zhang2023perception]. Furthermore, processing 3D data for lane detection requires substantial computational resources, which is particularly critical in AD applications where low latency is imperative. Reducing the processing time of algorithms without compromising their performance is also a significant challenge in such real-time applications.
Considering the aforementioned challenges and the importance of vision sensor-based 3D lane detection in accurate traffic scene understanding and parsing, we have accumulated existing research contributions and outcomes in this survey. Furthermore, before beginning this survey, we have posed the following questions: 1) Do existing datasets possess the potential for 3D lane detection in complex visual scenes? 2) What are the inference speeds of current methods, and can these methods meet the real-time requirements of autonomous vehicles? 3) Can current methods effectively perform 3D lane detection in complex visual scenes containing uncertainties such as fog and rain?
@@ I-C Contributions
    A comprehensive introduction to 3D lane detection techniques, defining the generic pipeline and explaining each step individually. This encourages newcomers in the field to quickly grasp prior knowledge and contributions from previous research, especially in the context of AD. To the best of our knowledge, this is the first survey on camera-based 3D lane detection.
@@ I-D Review Methodology
The research works discussed in this survey were retrieved using different keywords such as 3D lane detection in autonomous vehicles, vision-based 3D lane detection, and learning-based 3D lane detection. Most of the retrieved papers are directly relevant to the research topic, but there are also some exceptions, such as multi-modal methods [luo2022m, bai2018deep], and point cloud-based approaches [jung2018real], which are less related to the theme of this survey. Additionally, the aforementioned keywords were searched in multiple repositories, including Web of Science and Google Scholar, to ensure the retrieval of relevant content. Inclusion criteria ensured that a paper was recognized by experts in AD, based on factors such as citation count or the impact of prior work. It is worth mentioning that in the existing literature, no work on monocular 3D lane detection based on traditional methods was found. This may be because, unlike single-camera 2D lane detection, which only requires identifying pixels belonging to the lane in a 2D image, monocular 3D lane detection requires determining the 3D position information of the lane in the 3D space using the 2D image. Without the assistance of distance measuring sensors such as LiDAR or prediction through deep learning, it is difficult to achieve.
@@
The remainder of this paper is structured as follows: Section II provides an overview of related works for 3D lane detection in AD. Section III introduces performance evaluation of 3D lane detection. In section IV, commonly used datasets are introduced. Next, in section V, we introduce the current challenges and prospects for future 3D lane detection. Finally, Section VI concludes the paper.
@@
Figure 2: An overview of the 3D-LaneNet [garnett20193d] architecture. Information is processed in two parallel streams or pathways: the image-view pathway and the top-view pathway. This is called the dual-pathway backbone.